One professor grades homework by randomly choosing 3 out of 7 homework problems to grafe. If a student completed the selected problems, he or she will get a perfect score. Elise did 5 problems. What is the probability that she will get a perfect score.
Answer #1

Total selected problems by professor = 3

total assignments to be done by students = 7

total assignements Elise did =5

P( assignment completed by Elise will be chosen by professor in first attempt ) = 5/7

P(assignment completed by Elise will be chosen by professor in second attempt given completed assignment was chosen in first draw ) = 4/6

P(assignment completed by Elise will be chosen by professor in sthird attempt given completed assignment was chosen in first draw ) = 3/5

P( Elise get complete score) = probability that professor chose 3 assignments from those 5 assignements completed by Elise =   (5/7) * (4/6) * (3/5)

= 0.28571

So , there is 28.571 probability that Elise will get full grades